OpenStudy (anonymous):

Use the Pythagorean theorem to find the length of one leg if the hypotenuse is 45 m and the other leg is 34 m.

OpenStudy (anonymous):

\[\sqrt{45^2-34^2}\]

OpenStudy (sriram):

one leg=sqrt(45^2-34^2) =29.47

OpenStudy (llort):

and that equals = 57.008771549569

OpenStudy (anonymous):

\[ 45^2 = 34^2 + b^2 \] \[ 2025 = 1102 + b^2 \] \[ 2025 - 1102 = b^2 \] \[ 923 = b^2 \] \[\sqrt{923} = b\]